find three consecutive odd numbers such that the difference of 5 times the third number and three times the first number is 54 .. Thank you
Answer by ewatrrr(24785)

three consecutive odd numbers: n, (n+2), (n+4)
5(n+4)-3n = 54
5n + 20 - 3n = 54
2n = 34
n = 17
three consecutive odd numbers: 17, 19, 21
Checking:
105 - 51 = 54 CHECKS!